OpenStudy (anonymous):

Use the square root property to solve the equation. t^2=98 please use this formular -If t is a positive #, and if x^2=t, then x=sqrt t or x=-qurt t

Directrix (directrix):

According to your instructions, t^2=98 would, under the square roots property, become t = Sqrt(98) or t = -Sqrt(98). The next step is to simplify the Sqrt(98) by writing 98 as the product of a perfect square and another number. @mom2be012509